# Policymaker meetings as heteroscedasticity shifters:

### v1

@@ 5.2 FOMC meetings and heteroscedasticity of monetary policy shocks
It is well documented in the high-frequency identification literaturethat announcements made after FOMC meetings immediately impact financial variables. This change in financial indicators is associated with unforeseen changes in monetary policy, also referred to as monetary policy shocks.121212 See, for example, Cochrane and Piazzesi (2002), Faust et al. (2004), Gürkaynak et al. (2005), Gertler and Karadi (2015), Ramey (2016), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022)
We focus exclusively on the dates of FOMC meeting announcements including meetings, telephone and video conferences, unscheduled meetings, and sequential day meetings. Our baseline instrumental variable indicates the number of meetings occurring in a given month, including telephone conferences and unscheduled meetings. [^13] An alternative instrumental variable is a dummy with one’s in the months when the meetings occur and zero otherwise. Results are available upon request. Our instrument remains agnostic about the sign and size of the policy change. What matters is the event itself. [^14] We construct series using the FRASER St. Louis Fed website https://fraser.stlouisfed.org/title/federal-open-market-committee-meeting-minutes-transcripts-documents-677?browse=1980s Figure 3 plots the monthly time series of our baseline instrument starting from 1965 till 2007. As one can see the number of the meetings in a given month varies from zero to four.
Figure 3: FOMC meetings
Notes: The blue bars indicate the number of FOMC meetings per month, including telephone conferences and unscheduled meetings
Most of the papers that use FOMC meetings for identification purposes focus their analysis on the sample starting from 1994 onwards. This is done due to the fact the FED releases no public announcements about monetary policy decisions prior to 1994. Some researches, like, Jayawickrema and Swanson (2021) extend the sample backward to 1988. As our identification assumption remains agnostic about the direction of the change, size, we can use the sample back to 1965.
@@ 5.3.1 Benchmark specification
Figure 4 displays the impulse response functions of six endogenous variables to the structural monetary policy shock. We estimate those responses using Heteroskedssticity-IV LP estimator in our benchmark specification with $p=12$ without a trend. The dotted lines represent 68% point-wise confidence intervals, while the dot-dashed lines are simultaneous sup-t bands. Although simultaneous confidence bands are wider relative to point-wise, they are extremely useful, when one, for example, is interested in comparing short-run versus medium-run versus long-run effects. [^15] See section 5.3.4 below for comparison with an alternative simultaneous bounds based on Bonferroni inequality. Recent contribution regarding simultaneous confidence intervals are Jordà (2009), Inoue and Kilian (2016), Montiel Olea and Plagborg-Møller (2019), Inoue and Kilian (2022),Arias et al. (2023).
@@
Figure 4: Identification of monetary policy shock in the US. IRF estimates with point-wise and simultaneous bands. Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
@@
As observed, monetary policy tightening leads to a significant drop in output, which remains significant from 10 to 60 months using point-wise confidence bands. When considering joint significance, this decline is significant from approximately 15 to 35 months, a duration notably longer than previously reported results by Inoue and Kilian (2016). The results on inflation are similar to Wald-joint confidence intervals considered in that paper under a stationary VAR specification. In both cases, the price puzzle disappears once we consider simultaneous confidence intervals. It is important to keep in mind that our procedure of simultaneous confidence intervals construction allows for unit roots without pre-tests, while Inoue and Kilian (2016) requires transforming the model into a stationary one which may cause a pretesting bias.
5.3.2 Sensitivity to number of lags and trends
@@
Our inference methodology allows for a flexible trend specification. Appendix Figures 7-9 compare the benchmark model results (blue lines) with a more flexible models (golden lines): $(k=4,p=12)$, $(k=0,p=24)$, and $(k=4,p=24)$. The polynomial trend can account for slow and gradual changes in the mean values of stationary variables and the growth rates of unit root variables.
@@
Either including a trend or additional lags (or both) results in shorter-lived responses of the real GDP and weaker effects of the monetary shock on inflation and commodity prices. These results show that long-run effects of the monetary policy are sensitive to model specification. One possible explanation is that the trend (or additional lags) absorbs the long-run variation in the real GDP which is otherwise be attributed to monetary policy shocks. As a result, the more flexible specifications only detect the impact of monetary policy shocks that are neutral in the long run. This result is consistent with the panel regression findings at 10 year horizons in Jordà et al. (2020): the monetary tightening shocks can be non-neutral in the long run, while monetary loosening shocks are typically short lived.
@@
In the current literature, it is a convention to exclude the time trend from the monetary SVARs computed in levels (see, for example, Uhlig, 2005; Antolín-Díaz and Rubio-Ramírez, 2018). To facilitate the comparison, we will report only the results for the benchmark specification $(k=0,p=12)$ in the subsequent sections.
@@ 5.3.3 Smoothed LP
Appendix Figure 10 presents similar impulse response functions with smoothed point-wise and simultaneous bands for the benchmark specification. Compared to the raw LP estimates, Figure 4, the smoothed IRF results in a stronger impact of the shock on GDP at the long horizons, but a weaker impact on inflation in the long horizons. Notice that the confidence bounds got tighter as a result of the smoothing. This tightening occurs as a result of efficiency gains from imposing restrictions on the LP coefficient under the assumption of a correct specification of the VAR(12) model.
The smoothed LP estimators are based on a minimum distance estimator that imposes the VAR(12) model on the coefficients $C_{h}$. The distance between the VAR(12) model and the raw estimates was weighed with the inverse of the asymptotic variance of the individual LP estimators. Such weights are not fully optimal and could be further improved by using the inverse variance-covariance matrix of LP estimators as in Theorem 5. In practice, however, the optimal weights require some regularization in small samples. We leave this for future research.
@@ 5.3.4 Comparison with the union (Bonferroni) bounds
There are alternative existing methods that allow simultaneous inference on LP estimates without knowledge of the joint covariance matrix using so-called union or Bonferroni bounds (Inoue et al., 2023). Such an approach assumes the worst-case dependence between the individual components of the IRF estimates. As discussed in Montiel Olea and Plagborg-Møller (2019), the simultaneous sup-t bounds will be tighter than the union bounds in most cases. Appendix Figure 11 shows IRFs with sup-t versus union bands. Similar to Montiel Olea and Plagborg-Møller (2019) we find that Bonferroni confidence bands are noticeably wider relative to sup-t confidence bands.
5.3.5 Comparison with Cholesky
@@
Figure 5 compares the IRFs to a monetary policy shock identified Heteroscedasticity-IV LP to IRFs from Christiano et al. (1999). Both identification schemes use the smoothed LP IRF. The only difference is the estimation of the impact vector. As can be seen, the results of the new identification approach are very close to those of Christiano et al. (1999), which is a standard Cholesky identification scheme with a priori zeros impacts on real GDP, GDP deflator, and commodity price index. Specifically, the $\sup t$ test for the equality of the impact vectors for the two identification schemes given in Subsection 3.4 has $\sup t=1.2637$ with p-value $0.53$ (critical value at 32% is 1.8083). Without imposing restrictions, we obtain a similar conclusion.
@@
Notice that the confidence bounds for Heteroscedasticity-IV estimators account for uncertainty in the structural rotation matrix (which depends on both $\hat{\gamma}$ and $\hat{\Sigma}_{\eta}$), while the Cholesky approach only accounts for uncertainty in $\hat{\Sigma}_{\eta}$. As a result, the Cholesky IRF has misleadingly tighter confidence bounds that are only correct under the a priori zero restrictions on real GDP, GDP deflator and commodity prices implied by the identificaiton scheme.
@@
The classical paper Christiano et al. (1999) used recursive formulas for the computation of IRF. Appendix Figure 12 compares the results of smoothed LP Cholesky with the conventional VAR estimations under homoscedastic residuals. One can see that smoothed LP point estimates are only similar to the conventional VAR estimates at the short horizons. At the long horizons, the plots for real GDP, GDP deflator, and the commodity price index are substantially different. What is more surprising is that the smoothed LP point-wise confidence bounds are comparable and even tighter than the conventional VAR bootstrap bounds. It highlights the fact that the additional robustness of the proposed procedure does not automatically imply wider confidence bounds than its conventional alternative.
@@
Figure 5: Identification of monetary policy shock in the US. Comparison with Christiano et al. (1999). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and Cholesky-LP estimator (golden dot-dash) with corresponding 68% point-wise bands (blue dots and golden dashes).
@@ 5.3.6 Comparison with narrative sign restrictions
In the previous subsection, we found out that the zero restrictions on impacts on real GDP, GDP deflator, and the commodity price index cannot be rejected simultaneously. However, these restrictions were questioned by the literature. Uhlig (2005) replaced these zero restrictions with sign restrictions on impulse responses on short horizons. Such an approach by construction eliminates the price puzzle. In a more recent study by Antolín-Díaz and Rubio-Ramírez (2018) the authors further imposed restrictions on signs of monetary policy shock on particular dates to tighten the bounds. Both papers use Bayesian Gaussian homoscedastic SVAR models with uninformative priors over the structural rotation matrix for inference. [^16] Alternatively, Gafarov et al. (2018) considered frequentist bounds on partially identified sets of IRF using moment inequality methods. This approach results in wider confidence bounds since it is designed to cover the worst-case values of the unidentified structural rotations as opposed to the prior distribution weighted average.
Figure 6: Identification of monetary policy shock in the US. Comparison with narrative sign restrictions (Antolín-Díaz and Rubio-Ramírez, 2018). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and narrative sign restricted BVAR estimator (golden dot-dash) with corresponding 68% point-wise confidence and credibility bands (blue dots and golden dashes)
@@
There is an interesting discussion in the literature and some researchers argue that there is no price puzzle and this is a reality. For example, Mishkin (2007) argues that an increase in short-term interest rate increases the cost of building new houses and reduces housing activity. Through the supply cost channel that may cause an increase in prices in the short run. This example is from the construction sector, although similar discussions are shared in the agricultural sector. With monetary policy tightening, the cost of mortgages increases, which impacts production costs in the affected sectors.
@@ 6 Conclusion
This paper makes several contributions to the literature on SVARs with potential unit roots and co-integration. First, we propose a novel heteroscedasticity IV for monetary policy shocks that requires weaker assumtions than the conventional external IV approaches. As a result, we are able to construct much longer time series of IV. Second, we show how one can perform point-wise and simultaneous inference on IRF with such extenral IV using LP. We specifically generalize the HAC estimator of LP account for the non-statioanry regressors and show equivalence of dependent wild bootstrap with the analytical HAC inference. Third, we use the joint draws of the LP to construct minimum distance smoothed LP IRF estimator and the corresponding confidence bounds.
We then illustrate the usefullness of these procedures in two applications: a simulated data based on a realistic DSGE model and a monetary SVAR for the US. We find that the heterscedasticiy IV based on the number of FOMC meetings in a given month results in IRF that are very close to the classical Christiano et al. (1999) bounds and are different from a comparable BVAR with (narrative) sign restrictions that rule out price puzzle. In particular, our results suggest a much weaker long-run impact on inflation than the BVAR would imply.
@@ A.1 Sensitivity to model specification
Figure 7: Identification of monetary policy shock in the US. IRF estimates with a flexible trend ($k=4$ and $p=12$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=12$ (golden dot-dash line) with corresponding 68% point-wise bands
Figure 8: Identification of monetary policy shock in the US. IRF estimates with a additional lags ($k=0$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=0,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands
Figure 9: Identification of monetary policy shock in the US. IRF estimates with both flexible trend and additional lags ($k=4$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=24$ (golden dot-dash line) with corresponding 68% point-wise bands

### v2

@@ 5.2 FOMC meetings and heteroscedasticity of monetary policy shocks
It is well documented in the high-frequency identification literature that announcements made after FOMC meetings immediately impact financial variables. These changes in financial indicators are associated with unforeseen changes in monetary policy, commonly referred to as monetary policy shocks.161616 See, for example, Cochrane and Piazzesi (2002), Faust et al. (2004), Gürkaynak et al. (2005), Gertler and Karadi (2015), Ramey (2016), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022).
We focus exclusively on the dates of FOMC meeting announcements, including meetings, telephone and video conferences, unscheduled meetings, and sequential day meetings. Our baseline instrumental variable indicates the number of meetings that occur in a given month, including telephone conferences and unscheduled meetings. [^17] An alternative instrumental variable is a dummy with ones in the months when the meetings occur and zero otherwise. Results are available upon request. Our instrument remains agnostic about the sign and size of the policy change. What matters is the event itself. [^18] We construct series using the FRASER St. Louis Fed website https://fraser.stlouisfed.org/title/federal-open-market-committee-meeting-minutes-transcripts-documents-677?browse=1980s Figure 3 plots the monthly time series of our baseline instrument starting from 1965 till 2007. As can be seen, the number of meetings in a given month varies from zero to four.
Figure 3: Number of FOMC meetings per month from 1965 to 2007.
Notes: The blue bars indicate the number of FOMC meetings per month, including telephone conferences and unscheduled meetings.
Most of the papers that use FOMC meetings for identification purposes focus their analysis on the sample starting from 1994 onwards. This is because the FED released no public announcements about monetary policy decisions prior to 1994. Some scholars (for example, Swanson and Jayawickrema, 2023) extended the sample back to 1988. As our identification assumption is agnostic about the direction and size of the change, we can extend our sample back to 1965.
@@ 5.3.1 Benchmark specification
Figure 4 displays the impulse response functions of six endogenous variables to the structural monetary policy shock. We estimate those responses using the heteroscedasticity-IV LP estimator in our benchmark specification with $p=12$ and $k=0$ (without a trend). The dotted lines represent 68% point-wise confidence intervals, while the dashed lines are simultaneous sup-t bands. Although simultaneous confidence bands are wider relative to point-wise intervals, they are useful when, for example, one is interested in comparing short-run, medium-run, and long-run effects. [^19] See section 5.3.4 below for comparison with alternative simultaneous bounds based on the Bonferroni inequality. Recent contributions on simultaneous confidence intervals are Jordà (2009), Inoue and Kilian (2016), Montiel Olea and Plagborg-Møller (2019), Inoue and Kilian (2022), Arias et al. (2023).
Before we proceed to the estimation, it is worth noting that the hypothesis $\gamma_{1}=0$ is rejected in favor of the alternative $\gamma_{1}>0$ (Assumption 1) at 5% significance level. The corresponding test statistic is $t=1.77$, with a p-value equal to $0.039$ based on Algorithm 3. In other words, the variance of the Fed funds rate forecast errors is higher in the months with more FOMC meetings. This confirms earlier findings in the high-frequency identification literature that markets react to FOMC meeting announcements (see, for example, Faust et al., 2004).
@@
Figure 4: Identification of monetary policy shock in the U.S. IRF estimates with point-wise and simultaneous bands. The heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dash line).
@@
As evident from Figure 4, monetary policy tightening leads to a significant drop in output, which remains significant from 10 to 60 months using point-wise confidence bands. When considering joint significance, this decline is significant from approximately 15 to 35 months, a duration notably longer than previously reported results by Inoue and Kilian (2016). The results on inflation are similar to those obtained using Wald-joint confidence intervals considered in that paper under a stationary VAR specification. In both cases, the price puzzle disappears once we consider simultaneous confidence intervals. It is important to keep in mind that our procedure for constructing simultaneous confidence intervals allows for unit roots without pre-testing, while Inoue and Kilian (2016) requires transforming the model into a stationary one, which may cause a pre-testing bias.
5.3.2 Sensitivity to the number of lags and polynomial trends
@@
Our inference methodology allows for a flexible trend specification. Appendix Figures 7-9 compare the benchmark model results (blue lines) with more flexible models (golden lines): $(k=4,p=12)$, $(k=0,p=24)$, and $(k=4,p=24)$. The polynomial trend can account for slow and gradual changes in the mean values of stationary variables and the growth rates of unit root variables.
@@
Including either a trend or additional lags (or both) results in shorter-lived responses of real GDP and weaker effects of the monetary shock on inflation and commodity prices. These results show that the long-run effects of monetary policy are sensitive to model specification. One possible explanation is that the trend (or additional lags) absorbs the long-run variation in real GDP that is otherwise attributed to monetary policy shocks. As a result, the more flexible specifications only detect the impact of monetary policy shocks that are neutral in the long run. This result is consistent with the panel regression findings at 10-year horizons in Jordà et al. (2020): monetary tightening shocks can be non-neutral in the long run, while monetary loosening shocks are typically short-lived.
@@
In the current literature, it is conventional to exclude the time trend from monetary SVARs computed in levels (see, for example, Uhlig, 2005; Antolín-Díaz and Rubio-Ramirez, 2018). To facilitate the comparison, we will report only the results for the benchmark specification $(k=0,p=12)$ in the subsequent sections.
@@ 5.3.3 Smoothed LP
Appendix Figure 10 presents impulse response functions with smoothed point-wise and simultaneous bands for the benchmark specification. Compared to the raw LP estimates in Figure 4, the smoothed IRF results in a stronger impact of the shock on GDP at longer horizons but a weaker impact on inflation at longer horizons. Notice that the confidence bounds become tighter as a result of the smoothing. This tightening occurs as a result of efficiency gains from imposing restrictions on the LP coefficients under the assumption of a correct specification of the VAR(12) model.
The smoothed LP estimators are based on a minimum distance estimator that imposes the VAR(12) model on the coefficients $C_{h}$. The distance between the VAR(12) model and the raw estimates was weighted with the inverse of the asymptotic variance of the individual LP estimators. Such weights are not fully optimal and could be further improved by using the inverse variance-covariance matrix of LP estimators, as in Theorem 5. In practice, however, the optimal weights require some regularization in small samples. We leave this for future research.
@@ 5.3.4 Comparison with the union (Bonferroni) bounds
There are alternative existing methods that allow simultaneous inference on LP estimates without knowledge of the joint covariance matrix using so-called union or Bonferroni bounds (Inoue et al., 2023). Such an approach assumes the worst-case dependence between the individual components of the IRF estimates. As discussed in Montiel Olea and Plagborg-Møller (2019), the simultaneous sup-t bounds will be tighter than the union bounds in most cases. Appendix Figure 11 shows IRFs with sup-t versus union bands. Similar to Montiel Olea and Plagborg-Møller (2019), we find that Bonferroni confidence bands are noticeably wider relative to sup-t confidence bands.
5.3.5 Comparison with Cholesky-type identification
@@
Figure 5 compares the IRFs to a monetary policy shock identified using heteroscedasticity-IV LP to IRFs from Christiano et al. (1999). Both identification schemes use the smoothed LP IRF. The only difference is the estimation of the impact vector. As can be seen, the results of the new identification approach are very close to those of Christiano et al. (1999), which is a standard Cholesky identification scheme with a priori zero impacts on real GDP, GDP deflator, and commodity price index. Specifically, the $\sup t$ test for the equality of the impact vectors for the two identification schemes given in Subsection 3.4 has $\sup t=1.2637$ with a p-value $0.53$ (critical value at 32% level is 1.8083). Without imposing the apriori zero restrictions, we obtain a similar conclusion.
@@
Notice that the confidence bounds for the heteroscedasticity-IV estimators account for uncertainty in the structural rotation matrix (which depends on both $\hat{\gamma}$ and $\hat{\Sigma}_{\eta}$), while the Cholesky approach only accounts for uncertainty in $\hat{\Sigma}_{\eta}$. As a result, the Cholesky IRF has misleadingly tighter confidence bounds that are only correct under the a priori zero restrictions on real GDP, GDP deflator, and commodity prices implied by the identification scheme.
@@
The classical paper Christiano et al. (1999) uses recursive formulas for the computation of IRF. Appendix Figure 12 compares the results of the smoothed LP Cholesky method with conventional VAR estimations under homoscedastic residuals. One can see that smoothed LP point estimates are similar to the conventional VAR estimates only at the short horizons. At the long horizons, the plots for real GDP, GDP deflator, and the commodity price index are substantially different. What is more surprising is that the smoothed LP point-wise confidence bounds are comparable to, and even tighter than the conventional VAR bootstrap bounds. This highlights the fact that the additional robustness of the proposed procedure does not automatically imply wider confidence bounds than its conventional alternative.
@@
Figure 5: Identification of the monetary policy shock in the U.S.: Comparison with Christiano et al. (1999). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and Cholesky-LP estimator (golden dashed line) with corresponding 68% point-wise bands (blue dots and golden dashes).
@@ 5.3.6 Comparison with narrative sign restrictions
In the previous subsection, we found that the zero restrictions on impacts on real GDP, GDP deflator, and the commodity price index cannot be rejected simultaneously. However, these restrictions have been questioned in the literature. In particular, Uhlig (2005) replaced these zero restrictions with sign restrictions on impulse responses at short horizons. Such an approach eliminates the price puzzle by construction. In a more recent study, Antolín-Díaz and Rubio-Ramirez (2018) further imposed additional restrictions on the signs of the monetary policy shock at particular dates to tighten the bounds on IRF. Both papers use Bayesian Gaussian homoscedastic SVAR models with uninformative priors over the structural rotation matrix for inference. [^20] Alternatively, Gafarov et al. (2018) considered frequentist bounds on partially identified sets of IRF using moment inequality methods. This approach results in wider confidence bounds since it is designed to cover the worst-case values of the unidentified structural rotations as opposed to the prior distribution weighted average.
Figure 6: Identification of monetary policy shock in the U.S. Comparison with narrative sign restrictions (Antolín-Díaz and Rubio-Ramirez, 2018). Heteroscedasticity-IV LP estimator with $k=0,p=12$ (blue line) and narrative sign restricted BVAR estimator (golden dash) with corresponding 68% point-wise confidence and credibility bands (blue dots and golden dashes).
@@
There is an interesting discussion in the literature where some researchers argue that the price puzzle may reflects the reality. For example, Mishkin (2007) argues that an increase in the short-term interest rate raises the cost of building new houses and reduces housing activity. Through the supply cost channel this may cause an increase in prices in the short run. This example pertains to the construction sector. With monetary policy tightening, the cost of mortgages increases, which impacts production costs in the affected sectors.
@@ 6 Conclusion
This paper makes several contributions to the literature on SVARs with potential unit roots and cointegration. First, we propose a novel heteroscedasticity-baesd IV for monetary policy shocks that requires weaker assumptions than the conventional external IV approaches. As a result, we are able to construct much longer time series of instruments. Second, we show how one can perform point-wise and simultaneous inference on IRF with such heteroscedasticity-based external IV using the LP estimators. We specifically generalize the HAC estimator of LP account for the non-statioanry regressors and show equivalence of dependent wild bootstrap and the analytical HAC inference. Third, we use the joint draws of the LP estimates to construct minimum distance smoothed LP estimator and the corresponding confidence bounds.
We then illustrate the usefulness of these procedures in two applications: a simulated data based on a realistic DSGE model and a monetary SVAR for the U.S. We find that the heteroscedasticity IV based on the number of FOMC meetings in a given month results in IRF that are very close to the classical Christiano et al. (1999) bounds and are different from those obtained using a comparable BVAR with narrative sign restrictions that rule out the price puzzle. In particular, our results suggest a much weaker long-run impact on inflation than the BVAR would imply.
@@ A.1 Sensitivity to model specification
Figure 7: Identification of the monetary policy shock in the U.S. IRF estimates with a flexible trend ($k=4$ and $p=12$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=12$ (golden dashed line) with corresponding 68% point-wise bands.
Figure 8: Identification of the monetary policy shock in the U.S. IRF estimates with additional lags ($k=0$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=0,p=24$ (golden dashed line) with corresponding 68% point-wise bands.
Figure 9: Identification of the monetary policy shock in the U.S. IRF estimates with both a flexible trend and additional lags ($k=4$ and $p=24$). Heteroscedasticity-IV LP estimator $k=0,p=12$ (blue line) and $k=4,p=24$ (golden dashed line) with corresponding 68% point-wise bands.